Find coefficient of correlation, if variance of X series=2.25, standard deviation of Y series=4and regression equation of X on Y is X+0.3Y=1.8
1
Expert's answer
2020-03-09T09:36:22-0400

Sample regression coefficient x on y:

"\rho"xy=r*"\sigma"x/"\sigma"y

r - correlation coefficient between variables X and Y;

σx, σy - standard deviation calculated for the variables x b y

ρxy - coefficient in the equation

σy =4

σx= 1.5, if we extract the square root from the variance, we get the standard deviation

Transform the equation:

x+0.3y=1.8

0.3y=1.8-x, ρxy=0.3

Substitute in the first formula:

0.3=r*1.5/4

0.3=r*0.375

r=0.8

on a Cheddock scale

0.7 < r < 0.9: high.